an electron after being accelerated through potential difference of 10^4 V enters in a uniform magnetic field of 0.4 T perpendicular to its direction of motion. Calculate the radius of curvature of its trajectory.

Grade:10

2 Answers

Vikas TU
7 years ago
R = mv/qB
put all the values
mass of electron
q as charge of electron
B magnetic field
 
for velocity,
0.5mv^2 = qV
calcultae v and put in radius of curvature eqn.
laksh
5 years ago
as we know that force to move elecron in circular path will be balance by, mv2/r=qvb sin90
r=mv/qb
as mv=p
and p=(2meV)1/2
then r=(2meV)½/qb